A single conservative force acting on a particle within a system varies as ๐นโƒ— = (โˆ’๐ด๐‘ฅ + ๐ต๐‘ฅ

2

)๐‘–, ฬ‚

where A and B are constants, ๐นโƒ— is in newtons, and x is in meters. (a) Calculate the potentialย 

energy function U(x) associated with this force for the system, taking U=0 at x=0. Find (b)

the change in potential energy and (c) the change in kinetic energy of the system as theย 

particle moves from x=2m to x-3m.


1
Expert's answer
2021-05-26T12:55:12-0400

Gives

"F=(-Ax+Bx^2)i"

"U=-\intop F.dr"

Put Value


"U=-\smallint(-Ax+Bx^2)i.(dxi+dyj+dzk)"

"U(x)=\smallint(Ax-Bx^2)dx"

"U(x)=\frac{Ax^2}{2}-\frac{Bx^3}{3}+c"

Then

x=0,U=0

"U(x)=\frac{Ax^2}{2}-\frac{Bx^3}{3}"

"U(2)=\frac{A2^2}{2}-\frac{B2^3}{3}"

"U(2)=2A-2.67B"

U(3)="\frac{A3^2}{2}-\frac{B3^3}{3}"

U(3)=4.5A-9B

โˆ†U=4.5A -9B -(2A-2.67B)=2.5A -6.33B

โˆ†U=2.5A-6.33B

If consider particle alone as a system

The change in its kinetic energy is the work done by the force

On the particle

W=โˆ†K


For the entire system of which this particle of a mamber

This work is internal work and equal to the negative of the change in potential energy of the system

โˆ†K=-โˆ†U= -2.5A+6.33B